# The emergent dynamics of double-folded randomly branching ring polymers

### v1

@@ III.1 Discrete Double-Folded Ring Model
We consider discrete rings made of $N_{m}$ monomers and embedded in a face-centered cubic (FCC) lattice, with the mesh size (i.e., the space unit) set to $1$. Following previous works [3, 24] and the original elastic lattice polymer models [40, 41, 42, 43], we include the possibility that two successive monomers along the ring may occupy the same lattice site. In this case, the bond length between these monomers is 00 (referred to as the stored length), compared to $1$ for extended bonds (Fig. 1a). The double-folded nature of the model is then imposed by pairing each extended bond of the double-folded polymer with an opposing extended bond, as shown in Fig. 1a.
III.2 Handling the Underlying Lattice Tree
@@
A randomly double-folded ring polymer is characterized by an underlying randomly branched tree structure [1, 2, 3, 25, 6, 5, 26]. This tree is an undirected graph, whose nodes (vertices of the graph) are located on the occupied lattice sites and edges correspond to double-folded ring bonds. By definition, the tree is acyclic, and any pair of tree nodes is connected through a unique shortest path on the tree. Fig. 1 provides a visual $2D$ representation of a double-folded ring together with its underlying tree structure.
@@
In our previous work [24], we simulated double-folded ring polymers and characterized the corresponding tree conformations only during post-processing. This is unproblematic in systems with hard excluded volume interactions, where a ring configuration corresponds to a unique tree structure, because all spatially co-localized monomers belong to the same tree node (Fig. 1). However, in ideal double-folded rings, multiple tree nodes may occupy the same lattice site, introducing ambiguity in defining the underlying tree structure [24] and the statistical weight of ring conformations [44]
@@
To solve this problem, we extend the description of the double-folded ring by considering the degrees of freedom associated with the underlying tree. To this end, we define a ring conformation ${(\cal C)}$ by its set of ($N_{m}$) monomer positions in the embedding space, $\{\vec{r}_{1},\vec{r}_{2},...,\vec{r}_{N_{m}}\}$ together with its set of ($N_{n}$) tree node positions, $\{\vec{R}_{1},\vec{R}_{2},...,\vec{R}_{N_{n}}\}$, the ring and tree connectivity graphs, and information about which monomers belong to which tree node. For ideal systems we impose the constraint that nearest-neighbor tree nodes must not overlap, i.e. $|\vec{R_{I}}-\vec{R_{J}}|=1$, which is automatically fulfilled in systems with hard excluded volume interactions. As a consequence, there are small differences in the local structure of the present ideal rings and those of [24]. A summary of the data structure is presented in Table 2.
Ring Monomer $i$ Tree Node $I$
@@
In this context, the functionality $f$ of a tree node (the degree of a vertex) is defined by the number of tree edges coming out from the node. Specifically, a node with $f=1$ corresponds to a leaf or branch tip, $f=2$ to a linear section of the double-folded ring, and $f\geq 2$ to a branching point. For simplicity, and without loss of generality [5], here we limit the functionality to $f_{\text{max}}=3$.
III.3 Dynamic Monte Carlo Scheme for Mixed Dynamics
To investigate the dynamical properties of double-folded rings, we performed dynamic MC simulations. To this end, we used two types of monomer moves and one new type of tree node move (see first subsection below), together capturing the three possible types of local dynamics in double-folded polymers: reptons, creation/retraction of side branches, and Brownian motion of tree nodes. In addition, when considering volume exclusion (see systems studied in Section III.5), we employed a Metropolis algorithm, setting a very high interaction energy (i.e., greater than $30$ units of thermal energy) between nodes on the same lattice site, with no interaction otherwise.
@@ III.3.1 Trial moves
This occurs when a monomer hops longitudinally along the ring, crossing an extended bond to a neighboring tree node, thereby transporting a unit of stored length along the tree without changing its structure or the system’s energy. An example is shown in Fig. 2a.
@@
[…]
@@
Figure 2: (a)-(b): One monomer move involves hopping a randomly selected monomer into the nearest neighbor lattice site. (a) The Repton move; is responsible for the local redistribution of stored lengths. (b) The Hairpin move; is responsible for the creation or annihilation of side branches. The top row shows the double-folded ring, and the bottom row shows the underlying tree structure. (c)-(d) Two examples of tree node trial moves attempting to displace a randomly selected tree node to a randomly chosen site out of twelve possible nearest neighbors under a double-folding preserving constraint.
III.4 Mixed Dynamics and the Unit of Time
@@
A natural way to define a mixed dynamics is to randomly alternate monomer and tree node sweeps with probabilities $p_{m}\in[0,1]$ and $(1-p_{m})$ respectively. For $p_{m}=1$ this essentially generates the rings-in-an-array-of-obstacles dynamics from Ref. [24], while $p_{m}=0$ simulates the Brownian dynamics of quenched trees. Intermediate values mix the two types of motion in arbitrary proportions.
@@
    Perform $M_{s}$ trials, selecting a monomer move with probability $p_{m}N_{m}/M_{s}$ and a tree node move with probability $(1-p_{m})N_{n}/M_{s}$.
@@
In addition, we have introduced acceptance probabilities $0\leq q_{rep},q_{hp}\leq 1$ for the two monomer moves. For $q_{rep}=0$ and $q_{hp}=1$, all Repton moves are suppressed, resulting in dynamics that involve only the Hairpin moves. In the opposite case of $q_{rep}=1$ and $q_{hp}=0$, the only accepted monomer moves are Repton moves generating mass transport along the tree. While equal values $10\leq q_{rep}=q_{hp}<$ 1 generate a slowed down version of the original dynamics, other choices allow to arbitrarily manipulate the relative weight of Repton and Hairpin moves.
III.5 Systems Studied and Equilibrium Preparation
@@
We performed simulations for double-folded rings composed of $N_{m}=64,125,216,343,512,1000$ monomers on a three-dimensional FCC lattice with periodic boundary conditions for three types of systems: ideal rings, isolated self-avoiding rings, and rings in the melt state. For the ideal rings, there were no restrictions on the number of tree nodes per lattice site (i.e., no excluded volume interaction). For the self-avoiding rings and the rings in the melt state, we included excluded volume interactions (explained in Section III.3). Self-avoiding rings were simulated in a sufficiently large box, while a high lattice occupation number, $\phi\simeq 0.93$, was used for the melt state. A summary of the simulation parameters and equilibrium values for the studied systems is provided in Table I in Supplementary Material.
@@
All reported results were generated by first fully equilibrating the system according to the protocol of [24], using monomer moves for annealed double-folding rings (see Supplementary Material, Section I). Then, at the initial time of all simulation runs, an “equilibrated” state was randomly chosen.
IV Results
@@
To set the stage, we begin with a brief overview of the systems’ static properties in Section IV.1, which also serves to validate our simulation methods. In Sections IV.2 and IV.3, we analyze the emergent dynamics of monomers and CM for different combinations of trial moves.
IV.1 Static Properties
@@
Figure 3: Conformational statistics for isolated ideal rings, isolated self-avoiding rings, and rings in the melt state when subjected to Quenched Tree dynamics (QTree), ring in an array of obstacle-like (RiAO) dynamics, and the combination of the two (mixed). Inset panels show the local slopes of the data and the horizontal dashed lines correspond to the expectation scaling exponents for each system. The conformational statistics should be independent of the dynamics if the system condition remains the same and the results in the figure strongly confirm this assertion. Error bars are the same size or smaller than the symbols.
@@
In Figure 3 we present a number of results characterizing the secondary structure and the tertiary structure of our double folded rings along the lines of Ref. [21, 24]. Different colors distinguish results for ideal, self-avoiding and melt systems, while the symbols indicate the type of dynamics we used to acquire the data.
@@
Specifically we have calculated for all pairs $(i,j)$ of ring monomers their distance, $L(i,j)$, on the graph and their square distance, $R^{2}(i,j)$, in space. Averages as a function of the ring contour distance, $l=|i-j|\langle b\rangle$ are plotted in the top row of Fig. 3. The average bond length $\langle b\rangle$ is smaller than one, since the monomer bonds can be 00 or $1$ due to the presence of reptons in our model. Note that the data shown in the top row of Figure 3 is plotted as a function of the effective ring contour distance, $l_{\text{eff}}=l(1-l/N_{m})$, as a means of reducing finite ring size effects [21]. As a complement, we averaged $L(i,j)$ and $R^{2}(i,j)$ over all monomer pairs of our rings to calculate the average tree contour distance $\langle L\rangle$ and the mean-square gyration radius $\langle R_{g}^{2}\rangle$. Results as a function of the average number of tree nodes, $\langle N_{n}\rangle$, are plotted in the panels in the bottom row of Fig. 3. A summary of these values for the studied systems is provided in Table I in the Supplementary Material. Additionally, we calculated the effective local exponents, $\rho_{eff}$ and $\nu_{eff}$, from the slopes of the data points shown in the four panels of Fig. 3. The results are presented in the insets of the corresponding panels.
@@
In contrast, there are marked differences between the mean square spatial distance between ring monomers, $\langle R^{2}(l_{\text{eff}})\rangle$ and $\langle R^{2}(N_{n})\rangle$, for the different ring types. For ideal trees, this quantity is given by [45] $\langle R^{2}(l)\rangle=l_{K}\langle L(l)\rangle$, where the Kuhn length $l_{K}$ for the present model is simply given by the lattice constant. The fact that the independently calculated data sets for $\langle L(l_{\text{eff}})\rangle$ and $\langle R^{2}(l_{\text{eff}})\rangle$ look indistinguishable presents a simple and independent check of the path analysis on the tree. As expected, self-avoiding rings are more strongly swollen relative to ideal rings than rings in a melt. Again, our data are compatible with the expected crossover from local linear behavior with $\langle R^{2}(l)\rangle=l$ to $\langle R^{2}(l)\rangle\sim l^{2\nu}$ with $\nu=1/4$ for ideal trees, $\nu=1/2$ for self-avoiding trees, and $\nu=1/3$ for trees in the melt state. Note, however, the deviation of the effective exponents on small scales from the expected asymptotic values, which suggests the existence of corresponding deviations for effective dynamic exponents.
Importantly, for our present purposes, there are no noticeable differences between the data acquired through different dynamical schemes. Figure 3 thus validates our simulation methods and confirms that starting from an ensemble of equilibrated rings all ergodic dynamical schemes generate the same conformational statistics independently of whether the tree connectivity is randomly quenched or annealed.
IV.2 Overview of the Emergent Dynamics
Figure 4: Overview of emergent dynamics for different ring polymer systems: ideal (top row), isolated self-avoiding (middle row), and double-folded ring polymers in melts (bottom row). Results are shown for: monomer MSD, $g_{1}(t)$ (left column), CM MSD , $g_{3}(t)$, (middle column) and the size dependence of the asymptotic CM diffusion constants, $D(N_{n})$, (right column). Dynamics considered include: hairpin move, slithering dynamics, ring-in-an-array-of-obstacles (RiAO)-like dynamics, quenched tree dynamics (QTree), and a mix of RiAO and QTree dynamics. Insets show the local slopes of the data, with dashed lines indicating the expected exponents for QTree and RiAO dynamics (summarized in Table 1). Error bars are smaller than or comparable to the symbol sizes.
@@ IV.2 Overview of the Emergent Dynamics
Results for the monomer MSD, $g_{1}(t)$, the CM MSD, $g_{3}(t)$, and the size dependence of the asymptotic CM diffusion constants $D=\lim_{t\to\infty}\frac{g_{3}(t)}{6t}$, are presented in separate panels in the three columns of the figure. In addition, we report results obtained for the combination of repton and tree node moves in Fig. 9 in Appendix B.
While hairpin moves can locally redistribute accumulations of stored length, the overall dynamics emerging from their exclusive application is close to negligible: the final plateau is reached after approximately $\sim 30M_{s}$, with monomer and CM MSDs remaining strictly proportional to each other, i.e., $g_{1}(t)=N_{m}g_{3}(t)$. In contrast, the slithering dynamics emerging from the exclusive application of repton moves allows ring monomers to explore, in the long run, all sections of a wrapped, fixed tree. The confinement of the motion becomes visible notably earlier for the ring’s CM than for individual monomers. Importantly, it is only the combination of the two monomer moves that renders the RiAO dynamics ergodic. Interestingly, the absolute magnitude of the latter varies only weakly between the three considered systems.
@@
The situation is quite different in the melt case, where the hard-excluded volume interactions and the high occupation number ($\phi\sim 0.93$) of the lattice cells induce a strong reduction in the acceptance probability of hairpin and tree node moves. The early-time monomer dynamics, in particular, is now significantly slower for QTree dynamics than for RiAO dynamics. However, the anomalous diffusion in the intermediate regime appears to be faster for the QTree dynamics. Our data for the size dependence of the diffusion constant, $D(N_{n})$, (bottom right panel in Fig. 4) suggest that the Brownian QTree dynamics is, at least for the present melts, asymptotically faster than the RiAO dynamics.
@@
For ideal and isolated self-avoiding rings, the MSD resulting from the combination of ring monomer and tree node moves is unremarkable, displaying an intermediate behavior between the MSD emerging from the exclusive application of either monomer moves or node moves. Namely, with our definition of a MC sweep as the unit of time (see Section III.4 ), the mixing of 50% tree node moves and 50% of the much less effective monomer moves essentially slows down the motion by a factor of two in time. By contrast, and notably, something interesting happens in the melt case: the mixed dynamics is faster than either of the pure components, what we investigate in detail below.
IV.3 Scaling analysis of the Emergent Dynamics

### v2

@@ 3.1 Discrete Double-Folded Ring Model
We consider discrete rings made of $N_{m}$ monomers embedded in a face-centered cubic (FCC) lattice, with the mesh size (i.e., the space unit) set to $1$. Following previous works 3, 38 and the original elastic lattice polymer models 52, 53, 54, 55, we include the possibility that two successive monomers along the ring may occupy the same lattice site. In this case, the bond length between these monomers is 0 (referred to as stored length), compared to $1$ for extended bonds (Fig. 1a). The double-folded nature of the model is then imposed by pairing each extended bond of the double-folded polymer with an opposing extended bond, as shown in Fig. 1a.
3.2 Handling the Underlying Lattice Tree
@@
A randomly double-folded ring polymer is characterized by an underlying randomly branched tree structure 1, 2, 3, 23, 6, 5, 39. This tree is an undirected graph, whose nodes (vertices of the graph) are located on the occupied lattice sites and edges correspond to double-folded ring bonds. By definition, the tree is acyclic, and any pair of tree nodes is connected through a unique shortest path on the tree. Fig. 1 provides a visual $2D$ representation of a double-folded ring together with its underlying tree structure.
@@
In our previous work 38, we simulated double-folded ring polymers and characterized the corresponding tree conformations only during post-processing. This is unproblematic in systems with hard excluded volume interactions, where a ring configuration corresponds to a unique tree structure, because all spatially co-localized monomers belong to the same tree node (Fig. 1). However, in ideal double-folded rings, multiple tree nodes may occupy the same lattice site, introducing ambiguity in defining the underlying tree structure 38 and the statistical weight of ring conformations 56
@@
To solve this problem, we extend the description of the double-folded ring by considering the degrees of freedom associated with the underlying tree. To this end, we define a ring conformation $(\cal C)$} by its set of ($N_{m}$) monomer positions in the embedding space, $\{\vec{r}_{1},\vec{r}_{2},...,\vec{r}_{N_{m}}\}$, together with its set of ($N_{n}$) tree node positions, $\{\vec{R}_{1},\vec{R}_{2},...,\vec{R}_{N_{n}}\}$, the ring and tree connectivity graphs, and information about which monomers belong to which tree node. For ideal systems we impose the constraint that nearest-neighbor tree nodes must not overlap, i.e. $|\vec{R_{I}}-\vec{R_{J}}|=1$, which is automatically fulfilled in systems with hard excluded volume interactions. As a consequence, there are small differences in the local structure of the present ideal rings and those of 38. A summary of the data structure is presented in Table 2.
Ring Monomer $i$ Tree Node II
@@
In this context, the functionality $f$ of a tree node (the degree of a vertex) is defined by the number of tree edges coming out from the node. Specifically, a node with $f=1$ corresponds to a leaf or branch tip, $f=2$ to a linear section of the double-folded ring, and $f>2$ to a branching point. For simplicity, and without loss of generality 5, we limit the functionality to $f_{\text{max}}=3$ in this work (For a more detailed analysis of this choice, please see Sec. II in the supplementary material.).
3.3 Dynamic Monte Carlo Scheme for Mixed Dynamics
To investigate the dynamical properties of double-folded rings, we performed dynamic MC simulations. To this end, we used two types of monomer moves and one new type of tree node move (see first subsection below), together capturing the three possible types of local dynamics in double-folded polymers: reptons, creation/retraction of side branches, and Brownian motion of tree nodes. In addition, when considering volume exclusion (see systems studied in Section 3.5), we employed a Metropolis algorithm, setting a very high interaction energy (i.e., greater than $30$ units of thermal energy) between nodes on the same lattice site, with no interaction otherwise.
@@ 3.3.1 Trial moves
This occurs when a monomer hops longitudinally along the ring, crossing an extended bond to a neighboring tree node, thereby transporting a unit of stored length along the tree without changing its structure or the system’s energy. An example is shown in Fig. 2a. A detailed analysis of the repton distribution is provided in the Supplementary Material, Sec. III.
@@
This occurs when a monomer flanked by two zero-length bonds hops in a transverse direction, extending or creating a side branch. During the move, both zero-length bonds unfold, generating two new paired, oppositely oriented extended bonds. Consequently, the moved monomer is reassigned to a new node of functionality $f=1$. This new node connects to the node to which the monomer was originally associated, increasing the latter’s functionality by one, as well as the total number of tree nodes, $N_{n}$. Consequently, $N_{n}$ varies during a simulation involving hairpin moves. The corresponding inverse move removes an extended bond pair, leading to the shortening or complete removal (annihilation) of a side branch. An example is shown in Fig. 2b.
@@
Figure 2: (a)-(b):A monomer move involves hopping a randomly selected monomer into the nearest neighbor lattice site. (a) The Repton move transports stored length along the ring and locally redistributes it without altering the tree structure. (b) The Hairpin move is responsible for the creation or annihilation of side branches. The top row shows the double-folded ring, and the bottom row shows the underlying tree structure. (c)-(d) Two examples of tree node trial moves, where a randomly selected tree node attempts to move to one of the twelve nearest neighbors of the face-centered cubic (FCC) lattice used in our simulations, subject to a double-folding preserving constraint. For clarity, a two-dimensional schematic is shown here, while the actual simulations are performed in three dimensions.
3.4 Mixed Dynamics and the Unit of Time
@@
A natural way to define a mixed dynamics is to randomly alternate monomer and tree node sweeps with probabilities $p_{m}\in[0,1]$ and $(1-p_{m})$ respectively. For $p_{m}=1$ this essentially generates the rings-in-an-array-of-obstacles dynamics from Ref. 38, while $p_{m}=0$ simulates the Brownian dynamics of quenched trees. Intermediate values mix the two types of motion in arbitrary proportions.
@@
    Perform $M_{s}$ trials. For each trial, randomly select a move type: a monomer move with probability $p_{m}N_{m}/M_{s}$, or a tree node move with probability $(1-p_{m})N_{n}/M_{s}$. Since the move type is drawn independently for each trial, this procedure amounts to sampling with replacement. The number of monomer and node moves fluctuates from sweep to sweep, but their statistical proportions are satisfied on average.
@@
In addition, we have introduced acceptance probabilities $0\leq q_{rep},q_{hp}\leq 1$ for the two monomer moves. For $q_{rep}=0$ and $q_{hp}=1$, all Repton moves are suppressed, resulting in a dynamics that involve only the Hairpin moves. In the opposite case of $q_{rep}=1$ and $q_{hp}=0$, the only accepted monomer moves are Repton moves generating mass transport along the tree. While equal values $0\leq q_{rep}=q_{hp}<1$ generate a slowed down version of the original dynamics, other choices allow to arbitrarily manipulate the relative weight of Repton and Hairpin moves.
3.5 Systems Studied and Equilibrium Preparation
@@
We performed simulations for double-folded rings composed of $N_{m}=64,125,216,343,512,1000$ monomers on a three-dimensional FCC lattice with periodic boundary conditions for three types of systems: ideal rings, isolated self-avoiding rings, and rings in the melt state. For the ideal rings, there were no restrictions on the number of tree nodes per lattice site (i.e., no excluded volume interaction). For the self-avoiding rings and the rings in the melt state, we included excluded volume interactions (explained in Section 3.3). Self-avoiding rings were simulated in a sufficiently large box, while a high lattice occupation number, $\phi\simeq 0.93$, was used for the melt state. A summary of the simulation parameters and equilibrium values for the studied systems is provided in Table I in Supplementary Material.
@@
All reported results were generated by first fully equilibrating the system according to the protocol of 38, using monomer moves for annealed double-folding rings (see Supplementary Material, Fig. S1 in Section I). Then, at the initial time of all simulation runs, an “equilibrated” state was randomly chosen.
IV.  RESULTS
@@
To set the stage, we begin with a brief overview of the systems’ static properties in Section 4.1, which also serves to validate our simulation methods. In Sections 4.2 and 4.3, we analyze the emergent dynamics of monomers and CM for different combinations of trial moves.
4.1 Static Properties
@@
Figure 3: Conformational statistics for isolated ideal rings, isolated self-avoiding rings, and rings in the melt state when subjected to Quenched Tree dynamics (QTree), ring in an array of obstacle-like (RiAO) dynamics, and the combination of the two (mixed, with $p_{m}=0.5$). Insets show the local slopes of the data; horizontal dashed lines correspond to the expectation scaling exponents for each system (summarized in Table 1). The conformational statistics should be independent of the dynamics if the system condition remains the same and the results in the figure strongly confirm this assertion. Error bars are the same size or smaller than the symbols.
@@
In Figure 3 we present a number of results characterizing the secondary structure and the tertiary structure of our double folded rings along the lines of Ref. 35, 38. Different colors distinguish results for ideal, self-avoiding and melt systems, while the symbols indicate the type of dynamics we used to acquire the data.
@@
Specifically we have calculated for all pairs $(i,j)$ of ring monomers their distance, $L(i,j)$, on the tree and their square distance, $R^{2}(i,j)$, in space. Averages as a function of the ring contour distance, $l=|i-j|\langle b\rangle$ are plotted in the top row of Fig. 3. The average bond length $\langle b\rangle$ is smaller than one, since the monomer bonds can be 0 or $1$ due to the presence of reptons in our model. Note that the data shown in the top row of Figure 3 is plotted as a function of the effective ring contour distance, $l_{\text{eff}}=l(1-{|i-j|}/N_{m})$, as a means of reducing finite ring size effects 35. As a complement, we averaged $L(i,j)$ and $R^{2}(i,j)$ over all monomer pairs of our rings to calculate the average tree contour distance $\langle L\rangle$ and the mean-square gyration radius $\langle R_{g}^{2}\rangle$. Results as a function of the average number of tree nodes, $\langle N_{n}\rangle$, are plotted in the panels in the bottom row of Fig. 3. A summary of these values for the studied systems is provided in Table I in the Supplementary Material. Additionally, we calculated the effective local exponents, $\rho_{eff}$ and $\nu_{eff}$, from the slopes of the data points shown in the four panels of Fig. 3. The results are presented in the insets of the corresponding panels.
@@
In contrast, there are marked differences between the mean square spatial distance between ring monomers, $\langle R^{2}(l_{\text{eff}})\rangle$ and $\langle R^{2}(N_{n})\rangle$, for the different ring types. For ideal trees, this quantity is given by 57 $\langle R^{2}(l)\rangle=l_{K}\langle L(l)\rangle$, where the Kuhn length $l_{K}$ for the present model is simply given by the lattice constant. The fact that the independently calculated data sets for $\langle L(l_{\text{eff}})\rangle$ and $\langle R^{2}(l_{\text{eff}})\rangle$ look indistinguishable presents a simple and independent check of the path analysis on the tree. As expected, self-avoiding rings are more strongly swollen relative to ideal rings than rings in a melt. Again, our data are compatible with the expected crossover from local linear behavior with $\langle R^{2}(l)\rangle=l$ to $\langle R^{2}(l)\rangle\sim l^{2\nu}$ with $\nu=1/4$ for ideal trees, $\nu=1/2$ for self-avoiding trees, and $\nu=1/3$ for trees in the melt state. Note, however, the deviation of the effective exponents on small scales from the expected asymptotic values, which suggests the existence of corresponding deviations for effective dynamic exponents.
Importantly, for the present purpose, there are no noticeable differences between the data acquired from different dynamical schemes. Figure 3 thus validates our simulation methods and confirms that starting from an ensemble of equilibrated rings all ergodic dynamical schemes generate the same conformational statistics independently of whether the tree connectivity is randomly quenched or annealed.
4.2 Overview of the Emergent Dynamics
@@
Results for the monomer MSD, $g_{1}(t)$, the CM MSD, $g_{3}(t)$, and the size dependence of the asymptotic CM diffusion constants $D=\lim_{t\to\infty}\frac{g_{3}(t)}{6t}$, are presented in separate panels in the three columns of the figure. In addition, we report results obtained for the combination of repton and tree node moves in Fig. S6 in sec. V in the supplementary material.
While hairpin moves can locally redistribute accumulations of stored length, the overall dynamics emerging from their exclusive application is close to negligible: the final plateau is reached after approximately $\sim 30M_{s}$, with monomer and CM MSDs remaining strictly proportional to each other, i.e., $g_{1}(t)=N_{m}g_{3}(t)$. In contrast, the slithering dynamics emerging from the exclusive application of repton moves allows ring monomers to explore, in the long run, all sections of a wrapped, fixed tree. Notably, the confinement of the motion becomes visible earlier for the ring’s CM than for individual monomers. Importantly, it is only the combination of the two monomer moves that makes the RiAO dynamics ergodic. Interestingly, the absolute magnitude of the latter varies only weakly between the three considered systems.
@@
The situation is quite different in the melt case, where the hard-excluded volume interactions and the high occupation number ($\phi\sim 0.93$) of the lattice cells induce a strong reduction in the acceptance probability of hairpin and tree node moves, thus affecting both RiAO and QTree dynamics. As the slithering component remains active, it dominates the early-time monomer motion in RiAO dynamics. Consequently, at short times RiAO dynamics exhibits faster monomer motion than QTree dynamics. However, in the intermediate-time regime QTree dynamics shows slightly larger effective exponents than RiAO (inset in bottom left panel in Fig. 4). Overall, for the present ring sizes RiAO is faster than QTree dynamics. However, our data for the size dependence of the diffusion constant, $D(N_{n})$, (bottom right panel in Fig. 4) suggest that QTree dynamics is asymptotically faster than RiAO dynamics.
@@
For ideal and isolated self-avoiding rings, the MSD resulting from the combination of ring monomer and tree node moves is unremarkable, displaying an intermediate behavior between the MSD emerging from the exclusive application of either monomer moves or node moves. Namely, with our definition of a MC sweep as the unit of time (see Section 3.4 ), the mixing of 50% tree node moves and 50% of the much less effective monomer moves essentially slows down the motion by a factor of two in time. By contrast, and notably, something interesting happens in the melt case: the mixed dynamics is faster than either of the pure components, what we investigate in detail below.
Figure 4: Overview of emergent dynamics for different ring polymer systems: ideal (top row), isolated self-avoiding (middle row), and double-folded ring polymers in melts (bottom row). Results are shown for: monomer MSD, $g_{1}(t)$ (left column), CM MSD , $g_{3}(t)$, (middle column) and the size dependence of the asymptotic CM diffusion constants, $D(N_{n})$, (right column). Dynamics considered include: hairpin move, slithering dynamics, ring-in-an-array-of-obstacles (RiAO)-like dynamics, quenched tree dynamics (QTree), and a mix of RiAO and QTree dynamics with $p_{m}=0.5$. Insets show the local slopes of the data, with dashed lines indicating the expected exponents for QTree and RiAO dynamics (summarized in Table 1). Error bars are smaller than or comparable to the symbol sizes.
4.3 Scaling analysis of the Emergent Dynamics